# On Identification of Dynamic Treatment Regimes with Proxies of Hidden Confounders

### v1

@@ 1 Introduction
Estimating single time point optimal treatment regimes and multiple time point dynamic treatment regimes has received much attention in a diverse array of fields, including biostatistics, computer science, and economics. When data are obtained from a randomized controlled trial, A/B test, or a sequential multiple assignment randomized trial (SMART), the assumptions required for identification and estimation of optimal rules/regimes can be met by design. Alternatively, when one wishes to estimate the optimal regime from observational data, unverifiable assumptions such as the no unmeasured confounding assumption (NUCA) is typically invoked in point exposure settings and its time-varying analog termed sequential randomization assumption (SRA) is likewise invoked in time-varying treatment settings (Robins (1986), Murphy (2003)). In this article, we do not assume that experimental data are available, e.g. SMARTs, nor do we impose SRA. Instead, we allow for confounding by hidden factors for which variables with special causal structure, called proxies, are available, which we leverage to identify an optimal dynamic treatment regime in an observational study. Our results are also relevant beyond observational studies, as even when treatment assignment is randomized, there often is noncompliance. In a SMART with noncompliance, randomization can be viewed either as a valid instrumental variable, or alternatively as treatment confounding proxy, one of the two types of proxies we plan to leverage for identification.
@@ 1.1 Related Literature
There are several strands of research related to our work. The proximal causal inference framework, originally introduced in Miao et al. (2018) for the point exposure scenario, has been extended in many directions. Notably, Qi et al. (2023) and Shen and Cui (2022) solve the point exposure optimal treatment regime problem under unmeasured confounding using proxies. Proximal inference tools have also been applied to more complex regimes, such as mediation analysis, longitudinal studies, and off-policy evaluation in Tchetgen Tchetgen et al. (2024), Ying et al. (2023), Dukes et al. (2023), Singh (2021), Bennett and Kallus (2021), Miao et al. (2022), respectively. These works, similar to ours, require solving nested integral equations. In contrast, none require the challenging task of identifying the joint distribution of counterfactuals as a crucial step towards identifying an optimal dynamic regime, as Ying et al. (2023) only considers models for marginal counterfactual distributions indexed by static regimes, and Bennett and Kallus (2021) and Miao et al. (2022) study partially observed markov decision processes.
@@
One recent work that achieves identification of optimal dynamic treatment regimes under unmeasured confounding is Han (2021), using instrumental variables in addition to structural assumptions grounded in econometrics. Aside from Han (2021), to the best of our knowledge, there do not exist general identification results for optimal dynamic treatment regimes in the presence of unmeasured confounding. We fill this gap by leveraging proxy variables under proxy relevance conditions formally known as completeness conditions, with similar characteristics as those in Ying et al. (2023).
@@ 2.1 Notation
To simplify the exposition, we consider a two time point setting, though the results can be extended to any finite number of time points as shown in the Appendix. We suppose the data contains $n$ iid draws of a random vector $(U_{0},Y_{0},Z_{1},A_{1},W_{1},U_{1},Y_{1},Z_{2},A_{2},W_{2},Y_{2})$. Here, the $A$ variables are treatments, $U_{0}$ and $U_{1}$ represent time-varying unmeasured confounders of the effects of $(A_{1},A_{2})$, $Y_{0}$ is a baseline measurement of the outcome which may also confound the effect of subsequent treatments, $Y_{1}$ and $Y_{2}$ are post-treatment outcome measurements, and $Z$ and $W$ are proxies, whose causal structure is introduced in the next section. The overarching goal is to identify the optimal treatment regime based on past treatments and outcomes. In the two time point setting, we consider regimes at time point 2, ($d_{2}=d_{2}(y_{1},a_{1},y_{0})$) that depend on past outcomes, and treatment; and at time point 1 ($d_{1}=d_{1}(y_{0})$) that depend on the pretreatment outcome measurement. We note that allowing the $Y$ outcome process to be multivariate poses no additional theoretical difficulty, though the proxy conditions we introduce later would need to apply to these multivariate outcomes. We focus on the univariate outcome case to simplify the exposition. We formalize the framework in Section 3. We use the summation notation $\sum$, which in a slight abuse of notation, can also be interpreted as integrals for continuous variables. Also, we adopt the convention that uppercase variables denote random variables and lower case denote realizations of a random variable. At times, we write $f(a)$ instead of $f(A=a)$ to represent the density or probability mass function of random variable $A$ evaluated at $a$, but the abbreviation will be clear from context. E denotes the expectation operator. Finally, we adopt the $\overline{A}_{k}$ notation to represent a random variable and all its predecessors, i.e. $\overline{A}_{k}=(A_{1},\ldots,A_{k})$ or $\overline{A}_{k}=(A_{0},A_{1},\ldots,A_{k})$.
@@ 2.2 Assumptions
We next introduce formal causal assumptions underlying the data. First, we assume potential outcomes $Y_{2}(a_{1},a_{2})$ and $Y_{1}(a_{1})$ are well-defined and there is no interference, such that the former represents a unit’s outcome had, possibly contrary to fact, its treatment history been set to $(a_{1},a_{2})$; the latter is likewise defined. We also make the standard consistency assumption, so that $Y_{2}=Y_{2}(A_{1},A_{2})$ and $Y_{1}=Y_{1}(A_{1})$. Key to the approach is an assumption that the unobserved factors $U_{0},U_{1}$ would in principle suffice to account for time-varying confounding which we formalize as follows.
@@
$\{Y_{2}\left(a_{1},a_{2}\right),\overline{W}_{2}\}\perp\!\!\!\perp\{A_{2},\overline{Z}_{2}\}|\left(\overline{U}_{1},\overline{Y}_{1},A_{1}=a_{1}\right)$
$\left\{{Y}_{1}\left(a_{1}\right),Y_{2}\left(a_{1},a_{2}\right),W_{1}\right\}\perp\!\!\!\perp\{A_{1},Z_{1}\}|\left(U_{0},Y_{0}\right)$
$\{\overline{W}_{2},Y_{1}\}\perp\!\!\!\perp Z_{1}\mid A_{1},U_{0},Y_{0}.$
The first two lines of the assumption correspond to the assumptions from Ying et al. (2023); specifically Equations 2-4. The third assumption is an additional assumption that requires that $Z_{1}$ has no direct effect on $W_{2}$. The canonical situation where these conditions might be expected to hold would be when $W_{1}$ and $W_{2}$ are negative control outcomes and $Z_{1}$ and $Z_{2}$ are negative control exposures. In other words, $A_{1}$ and $A_{2}$ do not have direct effects on $W_{1}$ and $W_{2}$, respectively and $Z_{1}$ and $Z_{2}$ do not have direct effects on $Y_{1}$ and $Y_{2}$, respectively. In addition, $Z$ and $W$ should not affect each other. Note that Assumption 1 does not preclude $A_{1}$ from having an effect on $W_{2}$. A DAG that satisfies Assumption 1 is depicted in Figure 1 below. This DAG is similar to the one depicted in Figure 2 of Ying et al. (2023), with the addition of a pre-treatment confounder $Y_{0}$ and an intermediate outcome $Y_{1}$. The $Y$ in that DAG corresponds to $Y_{2}$ in Figure 1. The $Y_{0},U_{0}$ and $Y_{1},U_{1}$ are grouped together in a single node for increased readability, and represents that they have the same parent and child nodes. Previously, Tchetgen Tchetgen et al. (2024) and Ying et al. (2023) reanalyzed a study of the effect of an anti-rheumatic therapy called Methotrexate (MTX) on average number of tender joints at end of follow-up among patients with rheumatoid arthritis. They utilized erythrocyte sedimentation rate as a negative control exposure and patient’s global assessment as a negative control outcome. More generally, good candidates for negative control exposures include randomization in a SMART with noncompliance, or a time-varying instrumental variable in a longitudinal observational study. For more examples of both types of proxy variables, we refer the reader to Shi et al. (2020b).
@@
Figure 1: An example of a DAG where Assumption 1 holds.
@@ 3.1 Characterizing the optimal regime
We wish to solve the following optimal treatment regime problem. In particular, the goal is to identify $d_{2}^{opt}(\overline{Y}_{1},\overline{A}_{1})$ and $d_{1}^{opt}(Y_{0})$ that solve
$d_{2}^{opt}(\overline{Y}_{1},\overline{A}_{1}),d_{1}^{opt}(Y_{0})\in\underset{d_{1},d_{2}}{\arg\max}\ {E}\left\{Y_{2}\left(d_{1}(Y_{0}),d_{2}(Y_{1},A_{1},Y_{0})\right)\right\}.$
In other words, we seek to find the regime at time point 1 based on a baseline outcome $Y_{0}$ and regime at time point 2 based on a baseline outcome measurement $Y_{0}$, previous treatment $A_{1}$, and previous outcome $Y_{1}$ that will maximize the final outcome $Y_{2}$. We refer to ${E}\left\{Y_{2}\left(d_{1}(Y_{0}),d_{2}(Y_{1},A_{1},Y_{0})\right)\right\}$ as the value of a regime $(d_{1},d_{2})$. The regret of an estimated regime is the value of the optimal regime minus the value of the estimated regime. We point out that maximizing the final outcome is without loss of generality, as the intermediate potential outcomes $Y_{1}(a_{1})$ can be used in defining the value function (see Chapter 6.2.1 of Tsiatis et al. (2019) for a detailed discussion). It is well-known that $d_{1}^{opt}$ and $d_{2}^{opt}$ have the following equivalent and more convenient characterizations (e.g. Chapter 7.2 of Tsiatis et al. (2019) ):
@@
$V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right)$ $=\underset{a_{2}}{\max}\ {E}\left\{Y\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\right\}$
@@
Note that
${E}\left\{Y_{2}\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\right\}=\sum_{y_{2}}\frac{f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)}{f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)}y_{2},\ \text{ and }$
${E}\left\{V_{2}\left(a_{1},Y_{1}\left(a_{1}\right),Y_{0}\right)|Y_{0}=y_{0}\right\}=\sum_{y_{1}}V_{2}(a_{1},y_{1},y_{0})f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right),$
which clearly indicates that the quantities $f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)$ and $f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)$ are crucial for identification of the optimal regime. Note that the task of identifying optimal regimes in multiple time points is markedly distinct from and more difficult than identification of optimal regimes in the single time point case. This is due to the fact that to identify the optimal regime requires identification of the joint distribution of the counterfactuals $(Y_{2}(a_{1},a_{2}),Y_{1}(a_{1}))$. Previous work for observational longitidunal studies such as Michael et al. (2023) and Ying et al. (2023) provide identification results for functionals of the marginal distribution of the single counterfactual $Y_{2}(a_{1},a_{2})$ for some fixed $a_{1},a_{2}$, but not the joint $(Y_{2}(a_{1},a_{2}),Y_{1}(a_{1}))$. By the latent SRA, i.e. Assumption 1, the joint density of the counterfactuals can be written in terms of observable and the unmeasured confounders by a standard application of Robins’ g-formula. Unfortunately, without access to the unmeasured confounders, we must resort to alternative assumptions for identification.

### v2

# On Identification of Optimal Dynamic Treatment Regimes with Proxies of Hidden Confounders

@@ 1 Introduction
1.1 Background
Estimating single time point optimal treatment regimes and multiple time point dynamic treatment regimes has received much attention in a diverse array of fields, including biostatistics, computer science, and economics. When data are obtained from a randomized controlled trial, A/B test, or a sequential multiple assignment randomized trial (SMART), the assumptions required for identification and estimation of optimal rules/regimes can be met by design. Alternatively, when one wishes to estimate the optimal regime from observational data, unverifiable assumptions such as the no unmeasured confounding assumption (NUCA) are typically invoked in point exposure settings and its time-varying analog termed sequential randomization assumption (SRA) is likewise invoked in time-varying treatment settings (Robins (1986), Murphy (2003)). In this article, we do not assume that experimental data are available, e.g. SMARTs, nor do we impose SRA. Instead, we allow for confounding by hidden factors for which variables with special causal structure, called proxies, are available, which we leverage to identify an optimal dynamic treatment regime in an observational study. Our results are also relevant beyond observational studies, as even when treatment assignment is randomized, there often is noncompliance. In a SMART with noncompliance, randomization can be viewed either as a valid instrumental variable or a treatment confounding proxy, one of the two types of proxies we leverage for identification.
@@ 1.2 Related Literature
There are several strands of research related to our work. The proximal causal inference framework, originally introduced in Miao et al. (2018) for the point exposure scenario, has been extended in many directions. Notably, Qi et al. (2023) and Shen and Cui (2022) solve the point exposure optimal treatment regime problem under unmeasured confounding using proxies. Proximal inference tools have also been applied to more complex regimes, such as mediation analysis (Dukes et al., 2023), longitudinal studies (Tchetgen Tchetgen et al., 2024; Ying et al., 2023; Singh, 2021), and off-policy evaluation (Bennett and Kallus, 2021; Miao et al., 2022; Shi et al., 2022), respectively. These works, similar to ours, require solving nested integral equations. In contrast, none require the challenging task of identifying the joint distribution of counterfactuals as a crucial step towards identifying an optimal dynamic regime. Ying et al. (2023) only considers models for marginal counterfactual distributions indexed by static regimes. Bennett and Kallus (2021), Shi et al. (2022), and Miao et al. (2022) study off-policy evaluation with unmeasured confounding using proxies, but their approaches are not amenable to finding optimal policies/regimes because the nuisance bridge functions they introduce depend on the policy being evaluated.
@@
A recent work that achieves identification of optimal dynamic treatment regimes under unmeasured confounding is Han (2021), using instrumental variables in addition to structural assumptions grounded in econometrics. Shahn et al. (2022) estimates optimal regime structural nested mean models under unmeasured confounding using a parallel trends assumption. Besides Han (2021) and Shahn et al. (2022), to the best of our knowledge, there do not exist general identification results for optimal dynamic treatment regimes in the presence of unmeasured confounding. We fill this gap by leveraging proxy variables under proxy relevance conditions formally known as completeness conditions, with similar characteristics as those in Ying et al. (2023).
Finally, one work that achieves identification of the joint distribution of potential outcomes is that of Shpitser et al. (2023) (see Appendix B), though the authors did not explicitly mention this as their goal was identification of the marginal, nor did they make the connection to optimal treatment regimes. Shpitser et al. (2023) required solutions to similar integral equations as ours for identification, though our proof technique is markedly different. Moreover, beyond identification of the joint distribution of potential outcomes, we 1) provide closed forms of the necessary solutions to the integral equations in the fully discrete setting, 2) provide a generalization to more than 2 timepoints, and 3) in a setting where the downstream outcomes take on a finite number of values, we outline identification using a different type of bridge function.
@@ 2.1 Notation
To simplify the exposition, we consider a two time point setting, though the results can be extended to any finite number of time points as discussed in Section C of the supplementary material. We suppose the data contains $n$ iid draws of a random vector $(U_{0},Y_{0},Z_{1},A_{1},W_{1},U_{1},Y_{1},Z_{2},A_{2},W_{2},Y_{2})$. Here, the $A$ variables are treatments, $U_{0}$ and $U_{1}$ represent time-varying unmeasured confounders of the effects of $(A_{1},A_{2})$, $Y_{0}$ is a baseline measurement of the outcome which may also confound the effect of subsequent treatments, $Y_{1}$ and $Y_{2}$ are post-treatment outcome measurements, and $Z$ and $W$ are proxies, whose causal structure is introduced in the next section. The overarching goal is to identify the optimal treatment regime based on past treatments and outcomes. In the two time point setting, we consider regimes at time point 2, ($d_{2}=d_{2}(y_{1},a_{1},y_{0})$) that depend on past outcomes, and treatment; and at time point 1 ($d_{1}=d_{1}(y_{0})$) that depend on the pretreatment outcome measurement. We note that allowing the $Y$ outcome process to be multivariate poses no additional theoretical difficulty, though the proxy conditions we introduce later would need to apply to these multivariate outcomes. Moreover, in practice, there typically exist measured time-varying confounders. We can straightforwardly account for and utilize such variables, though we cannot identify optimal regimes based upon them, unless proxies of such variables are measured. We focus on the univariate outcome case without measured time-varying confounders solely to simplify the exposition. We formalize the framework in Section 3. We use the summation notation $\sum$, which in a slight abuse of notation, can also be interpreted as integral notation when random variables are continuous. Also, we adopt the convention that uppercase variables denote random variables and lower case denote realizations of a random variable. At times, we write $f(a)$ instead of $f(A=a)$ to represent the density or probability mass function of random variable $A$ evaluated at $a$, but the abbreviation will be clear from context. $E$ denotes the expectation operator. We adopt the $\overline{A}_{k}$ notation to represent a random variable and all its predecessors, i.e. $\overline{A}_{k}=(A_{1},\ldots,A_{k})$ or $\overline{A}_{k}=(A_{0},A_{1},\ldots,A_{k})$. We use $|A|$ to denote the cardinality of a discrete random variable $A$.
@@ 2.2 Assumptions
We next introduce formal causal assumptions underlying the data. First, we assume potential outcomes $Y_{2}(a_{1},a_{2})$ and $Y_{1}(a_{1})$ are well-defined and there is no interference, such that the former represents a unit’s outcome had, possibly contrary to fact, its treatment history been set to $(a_{1},a_{2})$; the latter is likewise defined. We also make the standard consistency assumption, so that $Y_{2}=Y_{2}(A_{1},A_{2})$ and $Y_{1}=Y_{1}(A_{1})$. Key to the approach is an assumption that the unobserved factors $U_{0},U_{1}$ suffice to account for time-varying confounding, and the absence of relationships between the proxies and other variables, which we formalize as follows.
@@
The following hold: $\{Y_{2}(a_{1},a_{2}),\overline{W}_{2}\}\perp\!\!\!\perp\{A_{2},\overline{Z}_{2}\}|(\overline{U}_{1},\overline{Y}_{1},A_{1}=a_{1})$, $\{{Y}_{1}(a_{1}),Y_{2}(a_{1},a_{2}),W_{1}\}\perp\!\!\!\perp\{A_{1},Z_{1}\}|(U_{0},Y_{0})$, and $\{\overline{W}_{2},Y_{1}\}\perp\!\!\!\perp Z_{1}\mid A_{1},U_{0},Y_{0}$.
The first two conditional independence statements correspond to the assumptions from Ying et al. (2023); specifically Equations 2-4. The third assumption is an additional assumption that requires that $Z_{1}$ has no direct effect on $W_{2}$. The canonical situation where these conditions might be expected to hold would be when $W_{1}$ and $W_{2}$ are negative control outcomes and $Z_{1}$ and $Z_{2}$ are negative control exposures. In other words, $A_{1}$ and $A_{2}$ do not have direct effects on $W_{1}$ and $W_{2}$, respectively and $Z_{1}$ and $Z_{2}$ do not have direct effects on $Y_{1}$ and $Y_{2}$, respectively. In addition, $Z$ and $W$ should not affect each other. Note that Assumption 1 does not preclude $A_{1}$ from having an effect on $W_{2}$. A Directed Acyclic Graph (DAG) that satisfies Assumption 1 is depicted in Figure 1 below. This DAG is similar to the one depicted in Figure 2 of Ying et al. (2023), with the addition of a pre-treatment confounder $Y_{0}$ and an intermediate outcome $Y_{1}$. The $Y_{0},U_{0}$ and $Y_{1},U_{1}$ are grouped together in a single node for increased readability, and represents that they have the same parent and child nodes. Previously, Tchetgen Tchetgen et al. (2024) and Ying et al. (2023) reanalyzed a study of the effect of an anti-rheumatic therapy called Methotrexate on average number of tender joints at end of follow-up among patients with rheumatoid arthritis. They utilized erythrocyte sedimentation rate as a negative control exposure and patient’s global assessment as a negative control outcome. More generally, good candidates for negative control exposures include randomization in a SMART with noncompliance, or a time-varying instrumental variable in a longitudinal observational study. For more examples of both types of proxy variables, we refer the reader to Shi et al. (2020b).
@@
Figure 1: An example of a DAG where Assumption 1 holds. The bold arrow emanating from $Y_{0}$ and $U_{0}$ represents arrows to every other node.
@@ 3.1 Characterizing the optimal regime
We wish to solve the optimal treatment regime problem, which requires identifying $d_{2}^{opt}(\overline{y}_{1},\overline{a}_{1})$ and $d_{1}^{opt}(y_{0})$ that solve
$d_{2}^{opt}(\overline{y}_{1},\overline{a}_{1}),d_{1}^{opt}(y_{0})\in\underset{d_{1},d_{2}}{\arg\max}\ {E}\left\{Y_{2}\left(d_{1}(Y_{0}),d_{2}(Y_{1},A_{1},Y_{0})\right)\right\}.$ (1)
In other words, we seek to find the regime at time point 1 based on a baseline outcome $Y_{0}$ and regime at time point 2 based on a baseline outcome measurement $Y_{0}$, previous treatment $A_{1}$, and previous outcome $Y_{1}$ that will maximize the final outcome $Y_{2}$. We refer to ${E}\left\{Y_{2}\left(d_{1}(Y_{0}),d_{2}(Y_{1},A_{1},Y_{0})\right)\right\}$ as the value of a regime $(d_{1},d_{2})$. The regret of an estimated regime is the value of the optimal regime minus the value of the estimated regime. We point out that maximizing the final outcome is without loss of generality, as the intermediate potential outcomes $Y_{1}(a_{1})$ can be used in defining the value function (see Chapter 6.2.1 of Tsiatis et al. (2019) for a detailed discussion). In principle, it is possible to find the optimal regime by identifying the value function for any fixed regime and solving the optimization from Equation 1 directly. However, this is typically not computationally feasible, since identification of the value function of a fixed regime using proxies requires estimating nuisance functions that themselves depend on the fixed regime (Bennett and Kallus, 2021; Shi et al., 2022; Miao et al., 2022). This implies that finding the optimal regime directly using Equation 1 would require estimating (on the order of) as many nuisance functions as there are policies, which could be infinite. We turn to the backwards induction approach to circumvent this issue. The backwards induction approach utilizes the well-known equivalent and more convenient characterizations of $d_{1}^{opt}$ and $d_{2}^{opt}$ (e.g. Chapter 7.2 of Tsiatis et al. (2019)):
@@
$V_{2}^{opt}\left(a_{1},\overline{y}_{1}\right)$ $=\underset{a_{2}}{\max}\ {E}\left\{Y_{2}\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\right\}$
@@
Notice that ${E}\left\{V_{2}^{opt}\left(a_{1},Y_{1}\left(a_{1}\right),Y_{0}\right)|Y_{0}=y_{0}\right\}=\sum_{y_{1}}V_{2}^{opt}(a_{1},y_{1},y_{0})f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)$ and ${E}\{Y_{2}\left(a_{1},a_{2}\right)|\overline{Y}_{1}\left(a_{1}\right)=\overline{y}_{1}\}=\sum_{y_{2}}f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)/f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)y_{2}$, which clearly indicates that the quantities $f\left(Y_{2}\left(a_{1},a_{2}\right)=y_{2},Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)$ and $f\left(Y_{1}\left(a_{1}\right)=y_{1}|y_{0}\right)$ are crucial for identification of the optimal regime. The task of identifying optimal regimes with multiple time points is markedly distinct from and more difficult than identification of optimal regimes in the single time point case. This is because identifying the optimal regime using Equation 2 requires identification of the joint distribution of the counterfactuals $\{Y_{2}(a_{1},a_{2}),Y_{1}(a_{1})\}$. Previous work for observational longitidunal studies such as Michael et al. (2023) and Ying et al. (2023) provide identification results for functionals of the marginal distribution of the single counterfactual $Y_{2}(a_{1},a_{2})$ for some fixed $a_{1},a_{2}$, but not the joint $\{Y_{2}(a_{1},a_{2}),Y_{1}(a_{1})\}$. By the latent SRA, i.e. Assumption 1, the joint density of the counterfactuals can be written in terms of observable and the unmeasured confounders by a standard application of Robins’ g-formula. Unfortunately, without access to the unmeasured confounders, we must resort to alternative assumptions for identification.